# Decoupling Continual Semantic Segmentation

### v1

@@ 1 Introduction
Continual Semantic Segmentation (CSS) addresses a practical scenario where new segmentation tasks with novel classes emerge over time [5, 13, 63]. A machine learning model must effectively learn these new classes while retaining previously learned old knowledge, ensuring that old knowledge is not forgotten. CSS has many real-world applications, such as autonomous driving [3], medical imaging [16, 47] and remote sensing [45]. As a dense prediction task, CSS is particularly susceptible to semantic shifts, where the model’s understanding of class relationships may evolve over time. Furthermore, similar to other continual learning tasks, CSS faces the challenge of catastrophic forgetting, where previously learned knowledge can be influenced and overwritten by new information.
@@
Several methods have been proposed to address the challenges in CSS. Data replay [30, 71] is effective but requires storing raw data/features, leading to linear memory growth and potential privacy concerns. Regularization techniques [4] penalize deviations in important model parameters for old classes, but often struggle to balance retention and plasticity. Pseudo-labeling [13, 44] and knowledge distillation [43, 48, 54] preserve old class knowledge by generating pseudo-labels or distilling knowledge from the old model, but still can cause background class drift when uncertain pixels are labeled as background. To mitigate this, the background weight transfer method [34, 60] initializes new class classifiers with background weights. However, it may mix new class features with background features, especially in cases of limited training data.
Most prior CSS methods employ a single-stage (encoder-decoder) segmentation paradigm, as seen in models such as DeepLab-V3 [9]. In these methods based on pixel-level multi-class classification, segmentation masks (intra-class consistency) and class labels (inter-class discrimination) are tightly coupled within a shared set of model parameters. This creates a significant challenge in the continual learning setting where the supervisory signals needed for inter-class discrimination between old and new categories are systematically removed, which partly causes catastrophic forgetting.
@@ 2 Related Work
Continual Semantic Segmentation The challenges in CSS include catastrophic forgetting, stability-plasticity dilemma, and semantic (background) shift [63]. PLOP [13] employs pseudo-labels generated by the prior model and multi-scale local distillation to preserve old knowledge. A series of works [6, 34, 40, 54, 67, 68] follow this way, aiming to enhance model stability by reducing information loss and preserving existing knowledge. For example, BalConpas [8] selectively distills the most relevant feature, and Cs2KCA [11] uses pixel-level features as a prototype for each class. Though Exemplar Replay [49, 52] has advanced significantly in CSS from three aspects: Sample Replay [7, 22, 31, 38, 61], Feature Replay [26, 42, 56, 59] and Auxiliary Data [29, 57], dynamic networks [1, 24, 46, 20, 50, 53, 55] stand out for their efficacy in preserving crucial parameters and enabling flexible task-specific adjustments [15], making them particularly advantageous in CSS. While recent CSS methods like CoMasTRe [15] and CoMFormer [6] based on Mask2Former [10] also explore decoupling strategies, with CoMasTRe focusing on objectness learning and classification separately, our framework follows a distinct detection-then-segmentation paradigm which shows superior performance.
Language-Driven Continual Learning Inspired by the observation that humans effectively acquire new visual knowledge through language and motivated by the broad applications of pre-trained vision-language models (VLMs) [18, 37], previous studies have explored VLMs in continual learning classification tasks [21, 65, 69]. However, applications of VLMs for Continual Semantic Segmentation (CSS) are much less explored, primarily because the dense annotations required for segmentation cannot be directly leveraged by these models; only a few studies have explored their potential in Weakly Supervised Continual Semantic Segmentation [58]. To our best knowledge, our work is the first to leverage VLMs in dense-annotated CSS.
@@ 3.2.1 Text Encoding
For each of the $c_{t}$ classes, a number of $M+1$ descriptive phrases are generated (see Section B.1 for the complete template and process). Among them, $M$ phrases are generated by the large language model [33] using a pre-defined prompt [2, 35]. Unlike conventional approaches that generate complete sentences with redundant linguistic elements, our method employs concise phrasal descriptions (e.g., adjective + class name) to capture essential visual attributes. This focused representation enhances feature distinctiveness and improves class discrimination capabilities. Each phrase is then sent to a pre-trained text encoder to obtain the corresponding text embedding. Consequently, for the $k$-th class ($k=1,2,\ldots,c_{t}$), a set of $M+1$ text embeddings $\{\mathbf{g}_{k,1},\mathbf{g}_{k,2},\ldots,\mathbf{g}_{k,M+1}\}$ are generated. The purpose of generating multiple versions of textual descriptions and corresponding text embeddings for each class is to increase diversity and enhance the robustness of the subsequent aggregation step.
@@ 3.2.2 Task-Specific LoRAs
In our framework, we initialize the pre-trained image encoder (with a Swin Transformer backbone) and cross-attention module from Grounding DINO [27] which are endowed with robust feature extraction and cross-modal alignment capabilities. However, while these pre-trained components provide reliable general representations, they are not specifically optimized for the nuanced requirements of learning novel classes incrementally.
Here, task-specific LoRA [17] adapters are added into both the image encoder and the cross-attention module to learn each new segmentation task. For the image encoder, these lightweight adapters are embedded in each self-attention layer, enhancing discriminative feature extraction for new classes while maintaining the output structure as $\mathbf{V}_{i}\in\mathbb{R}^{N\times d}$ for each input image $\mathbf{x}_{i}$, where $N$ represents visual tokens and $d$ is the embedding dimension. In the cross-attention module, which coordinates bidirectional information flow between modalities, LoRA adapters are incorporated in the linear projections that generate attention ‘keys’ and ‘values’. This approach enables effective text-guided visual feature optimization through multiple interaction layers, producing enhanced representations $\mathbf{V}^{\prime}_{i}\in\mathbb{R}^{N\times d}$ and $\mathbf{E}^{\prime}_{i}\in\mathbb{R}^{c_{t}\times d}$ that better capture cross-modal relationships while maintaining computational efficiency.
@@ 3.3.2 Prompt Generation (pGen)
Replay
MicroSeg-M [66] 82.9 60.1 77.5 82.0 47.3 73.3 78.9 59.2 70.1
SSUL-M [7] 77.38 22.43 74.76 79.3 55.1 73.5 78.8 49.7 71.9 75.3 54.1 65.2
RECALL [31] 67.9 53.5 68.4 66.6 50.9 64.0 65.7 47.8 62.7 59.5 46.7 54.8
SATS-M [36] 81.44 70.02 78.72 80.37 64.54 76.61 76.21 61.62 69.27
@@
Data-free
MiB* [4] 69.91 20.63 67.45 75.48 49.41 68.47 36.71 12.12 30.82 12.20 13.19 12.61
PLOP* [13] 74.15 35.58 72.04 75.49 49.66 69.34 64.09 20.12 53.11 44.03 15.51 30.45
@@
MicroSeg [66] 81.9 54.0 75.2 80.5 40.8 71.0 73.5 53.0 63.8
CoMFormer [6] 75.35 24.06 72.91 74.68 54.30 71.12 70.78 32.24 61.60
@@
IPSeg [60] 81.4 62.4 76.9 82.4 52.9 75.4 80.0 61.2 71.0
SATS* [36] 77.42 61.07 74.41 80.24 61.17 75.70 78.38 62.02 74.48 64.27 58.66 61.60
[…]
SSUL [7] 79.7 55.3 73.9 78.1 33.4 67.5 74.3 51.0 63.2
@@ 3.4 Class-Agnostic Segmentation and Aggregation
The Class-Agnostic Segmentation (CAS) module utilizes a segmentation foundation model, and we currently employ SAM (Segment Anything Model) [23] for this role. SAM works in a class-agnostic manner and can generate high-quality segmentation results for the input image solely based on spatial prompt conditions. Specifically, for each class-specific prompt $\mathbf{p}_{k}$ (from above section), SAM outputs a corresponding mask $\mathcal{M}_{k}$ and a confidence score. These individual class-specific masks are aggregated into a final semantic segmentation map.
It is worth noting that the CAS module is frozen in the whole continual learning process. However, when the model learns task $t$, the outputs from CAS are used to generate the segmentation loss function. In our framework, the same segmentation loss as that for SAM training is adopted to train the learnable class-specific prompt generators, and the loss is also used together with the asymmetric loss [39] to train the task-specific LoRAs in the LTCD module.
During model inference, given a test image, the task-specific and class-specific components of each task are plugged into the segmentation system to segment image regions specifically for those classes estimated appearing in the input image. Such inference process is run over all tasks and the results are collected for the final segmentation. The final segmentation is created by aggregating individual class masks. While pixel ownership is unambiguous for the vast majority of pixels, any rare spatial overlaps are resolved using a confidence-based approach. The pixel is assigned to the class with the highest confidence score from the SAM decoder. Notably, this process requires no task ID, with time linear complexity with respect to the number of learned tasks $T$. Since only one adapter resides in memory at any time, maximum memory usage remains identical to that in single-task models. For time-sensitive applications, memory-time tradeoff techniques [41] can be used to reduce inference latency.
@@
Data-free
MiB* [4] 39.02 16.73 31.29 36.68 9.81 27.77 34.22 5.26 24.29
PLOP* [13] 40.38 13.41 31.52 39.46 12.58 30.10 38.12 7.32 27.39
@@ 4.1 Experimental Setup
Datasets Following previous studies [4, 13], we evaluate our method on PASCAL VOC2012 [14] and ADE20K [70]. PASCAL VOC2012 contains 20 object classes while ADE20K presents a more challenging scenario with 150 classes. Dataset statistics are in Appendix C.
CSS Settings. The training process is divided into $T$ tasks with certain protocol. The overlapped protocol, which is more challenging and realistic compared to the disjoint setting, allows images to contain old classes previously learned and future classes to be learned [36]. For these images, annotations are provided only for the current task’s classes and image regions corresponding to both old and future classes are annotated as background. Our experiments use this overlapped setting throughout. For instance, in the VOC2012 10-1 (11 tasks) setting, the model first learns to segment 10 classes, then incrementally learns one new class in each of 10 new tasks.
Implementation details. The model was trained for 5 epochs on PASCAL VOC2012 and 20 epochs on ADE20K using AdamW [28] optimizer (initial learning rate $1\times 10^{-4}$, weight decay 0.05) with a polynomial learning rate schedule. For the LTCD module, we set the similarity threshold $\tau$ to 0.3 for PASCAL VOC2012 and 0.2 for ADE20K, ensuring an appropriate quantity of selected visual tokens in $\mathbf{V}_{i}^{\text{sel}}$, with $M=30$ descriptive phrases per class and LoRA rank $L=32$. The SPG module uses $m=6$ prompt tokens with a single hidden layer, and pGen input lengths of 512 and 1024 for PASCAL VOC2012 and ADE20K respectively, with the longer length for ADE20K accommodating its complex scene composition and higher object density. All experiments were conducted with batch size 24 on 8 NVIDIA 3090 GPUs.
@@ 4.2 Main Results
MiB [4] 3.27 6.35 5.91 7.08 6.72 6.80 39.50 20.07 24.70
PLOP [13] 24.96 5.38 8.18 15.41 6.55 8.67 28.54 15.95 18.95
MiB+NeST [51] 13.44 6.72 7.68 10.09 7.51 7.77 38.31 21.3 25.35
PLOP+NeST [51] 27.98 6.53 8.26 17.57 7.84 10.15 33.89 17.84 21.66
@@ 4.5 Additional Studies
Finally, we analyse the computational cost of our approach. Similar to previous continual learning studies which employ task-specific modules [25], our method introduces computational overhead during inference primarily from switching task-specific LoRA parameters. The inference time for a single task is 0.3 second per image, and the full process for six tasks in the (15-1) setting takes about 1.9 seconds per image. In comparison, most existing single encoder-decoder based CSS methods takes around 0.5 second per image. Despite leveraging foundation models, our parameter growth remains highly manageable. Each class requires only 8MB for the pGen module (0.225% of total model size), while all tasks’ LoRA adapters collectively need merely 22.18MB (0.62% of total model size). Each task requires relatively few trainable parameters per increment [19], yet achieves state-of-the-art performance with remarkable gains in challenging settings, demonstrating practical applicability comparable to offline-trained models. Given our method’s substantial performance gains and affordable storage overhead, this trade-off is well-justified for non-realtime applications.

### v2

@@ 1 Introduction
Continual Semantic Segmentation (CSS) addresses a practical scenario where new segmentation tasks with novel classes emerge over time [cermelliModelingBackgroundIncremental2020, douillardPLOPLearningForgetting2021, 2024SurveyCISS]. A machine learning model must effectively learn these new classes while retaining previously learned old knowledge, ensuring that old knowledge is not forgotten. CSS has many real-world applications, such as autonomous driving [Lidarss], medical imaging [mediccss, IncreNucleiSegmentation] and remote sensing [remotecss]. As a dense prediction task, CSS is particularly susceptible to semantic shifts, where the model’s understanding of class relationships may evolve over time. Furthermore, similar to other continual learning tasks, CSS faces the challenge of catastrophic forgetting, where previously learned knowledge can be influenced and overwritten by new information.
@@
Several methods have been proposed to address the challenges in CSS. Data replay [recall, 2023datareplay] is effective but requires storing raw data/features, leading to linear memory growth and potential privacy concerns. Regularization techniques [CVPR2020ModelingTB] penalize deviations in important model parameters for old classes, but often struggle to balance retention and plasticity. Pseudo-labeling [douillardPLOPLearningForgetting2021, pseudo2019] and knowledge distillation [su2024Distillation, wang2024Distillation, yang2023distillation] preserve old class knowledge by generating pseudo-labels or distilling knowledge from the old model, but still can cause background class drift when uncertain pixels are labeled as background. To mitigate this, the background weight transfer method [ECCV2024MitigatingBS, ipseg] initializes new class classifiers with background weights. However, it may mix new class features with background features, especially in cases of limited training data.
Most prior CSS methods employ a single-stage (encoder-decoder) segmentation paradigm, as seen in models such as DeepLab-V3 [chen2017rethinking]. In these methods based on pixel-level multi-class classification, segmentation masks (intra-class consistency) and class labels (inter-class discrimination) are tightly coupled within a shared set of model parameters. This creates a significant challenge in the continual learning setting where the supervisory signals needed for inter-class discrimination between old and new categories are systematically removed, which partly causes catastrophic forgetting.
@@ 2 Related Work
Continual Semantic Segmentation The challenges in CSS include catastrophic forgetting, stability-plasticity dilemma, and semantic (background) shift [2024SurveyCISS]. PLOP [douillardPLOPLearningForgetting2021] employs pseudo-labels generated by the prior model and multi-scale local distillation to preserve old knowledge. A series of works [cermelliCoMFormerContinualLearning2023, ECCV2024MitigatingBS, shang2023distillation, yang2023distillation, InheritwithDistill, zhao2023distillation] follow this way, aiming to enhance model stability by reducing information loss and preserving existing knowledge. For example, BalConpas [ECCV2024StrikeStrikeaBalanceinContinualPanopticSegmentation] selectively distills the most relevant feature, and Cs2KCA [ECCV2024Cs2KCA] uses pixel-level features as a prototype for each class. Though Exemplar Replay [CVPR2022CL, yan2021DER] has advanced significantly in CSS from three aspects: Sample Replay [samrep2021ssul, samrep2022pin, samrep2021RECALL, samrep2017iCaRL, samrep2024tikp], Feature Replay [ferep2022dynamic, ferep2022, ferep2022semi, ferep2020] and Auxiliary Data [AD2021, AD2023foundation], dynamic networks [DA2022decomposed, DA2022continual, DA2023fairness, DA2023effects, DA2023endpoints, DA2022deep, DA2022learning] stand out for their efficacy in preserving crucial parameters and enabling flexible task-specific adjustments [CoMasTRe2024], making them particularly advantageous in CSS. While recent CSS methods like CoMasTRe [CoMasTRe2024] and CoMFormer [cermelliCoMFormerContinualLearning2023] based on Mask2Former [Mask2Former] also explore decoupling strategies, with CoMasTRe focusing on objectness learning and classification separately, our framework follows a distinct detection-then-segmentation paradigm which shows superior performance.
Language-Driven Continual Learning Inspired by the observation that humans effectively acquire new visual knowledge through language and motivated by the broad applications of pre-trained vision-language models (VLMs) [Huang_Huang_Zhang_Tian_Feng_Zhang_Xie_Li_Zhang_2023, Radford_Kim_Hallacy_Ramesh_Goh_Agarwal_Sastry_Amanda_Mishkin_Clark_et_al._2021], previous studies have explored VLMs in continual learning classification tasks [Knowledge-Aware, wentaozhang2024continual, zheng2024large]. However, applications of VLMs for Continual Semantic Segmentation (CSS) are much less explored, primarily because the dense annotations required for segmentation cannot be directly leveraged by these models; only a few studies have explored their potential in Weakly Supervised Continual Semantic Segmentation [FoundationModelDWISS]. To our best knowledge, our work is the first to leverage VLMs in dense-annotated CSS.
@@ 3.2.1 Text Encoding
For each of the $c_{t}$ classes, a number of $M+1$ descriptive phrases are generated (see Section B.1 for the complete template and process). Among them, $M$ phrases are generated by the large language model [OpenAI_2023] using a pre-defined prompt [Bai_Xia_2023, Pratt_Liu_Farhadi_2022]. Unlike conventional approaches that generate complete sentences with redundant linguistic elements, our method employs concise phrasal descriptions (e.g., adjective + class name) to capture essential visual attributes. This focused representation enhances feature distinctiveness and improves class discrimination capabilities. Each phrase is then sent to a pre-trained text encoder to obtain the corresponding text embedding. Consequently, for the $k$-th class ($k=1,2,\ldots,c_{t}$), a set of $M+1$ text embeddings $\{\mathbf{g}_{k,1},\mathbf{g}_{k,2},\ldots,\mathbf{g}_{k,M+1}\}$ are generated. The purpose of generating multiple versions of textual descriptions and corresponding text embeddings for each class is to increase diversity and enhance the robustness of the subsequent aggregation step.
@@ 3.2.2 Task-Specific LoRAs
In our framework, we initialize the pre-trained image encoder (with a Swin Transformer backbone) and cross-attention module from Grounding DINO [liu2024groundingdinomarryingdino] which are endowed with robust feature extraction and cross-modal alignment capabilities. However, while these pre-trained components provide reliable general representations, they are not specifically optimized for the nuanced requirements of learning novel classes incrementally.
Here, task-specific LoRA [hu2021lora] adapters are added into both the image encoder and the cross-attention module to learn each new segmentation task. For the image encoder, these lightweight adapters are embedded in each self-attention layer, enhancing discriminative feature extraction for new classes while maintaining the output structure as $\mathbf{V}_{i}\in\mathbb{R}^{N\times d}$ for each input image $\mathbf{x}_{i}$, where $N$ represents visual tokens and $d$ is the embedding dimension. In the cross-attention module, which coordinates bidirectional information flow between modalities, LoRA adapters are incorporated in the linear projections that generate attention ‘keys’ and ‘values’. This approach enables effective text-guided visual feature optimization through multiple interaction layers, producing enhanced representations $\mathbf{V}^{\prime}_{i}\in\mathbb{R}^{N\times d}$ and $\mathbf{E}^{\prime}_{i}\in\mathbb{R}^{c_{t}\times d}$ that better capture cross-modal relationships while maintaining computational efficiency.
@@ 3.3.2 Prompt Generation (pGen)
Replay MicroSeg-M [microseg] 82.9 60.1 77.5 82.0 47.3 73.3 78.9 59.2 70.1
SSUL-M [samrep2021ssul] 77.38 22.43 74.76 79.3 55.1 73.5 78.8 49.7 71.9 75.3 54.1 65.2
RECALL [samrep2021RECALL] 67.9 53.5 68.4 66.6 50.9 64.0 65.7 47.8 62.7 59.5 46.7 54.8
SATS-M [SATS] 81.44 70.02 78.72 80.37 64.54 76.61 76.21 61.62 69.27
@@
Data-free MiB* [CVPR2020ModelingTB] 69.91 20.63 67.45 75.48 49.41 68.47 36.71 12.12 30.82 12.20 13.19 12.61
PLOP* [douillardPLOPLearningForgetting2021] 74.15 35.58 72.04 75.49 49.66 69.34 64.09 20.12 53.11 44.03 15.51 30.45
@@
MicroSeg [microseg] 81.9 54.0 75.2 80.5 40.8 71.0 73.5 53.0 63.8
CoMFormer [cermelliCoMFormerContinualLearning2023] 75.35 24.06 72.91 74.68 54.30 71.12 70.78 32.24 61.60
$\dagger [CoMasTRe$2024] 75.13 69.51 74.86 79.73 51.93 73.11 69.77 43.62 63.54
IPSeg [ipseg] 81.4 62.4 76.9 82.4 52.9 75.4 80.0 61.2 71.0
SATS* [SATS] 77.42 61.07 74.41 80.24 61.17 75.70 78.38 62.02 74.48 64.27 58.66 61.60
@@
SSUL [samrep2021ssul] 79.7 55.3 73.9 78.1 33.4 67.5 74.3 51.0 63.2
@@ 3.4 Class-Agnostic Segmentation and Aggregation
The Class-Agnostic Segmentation (CAS) module utilizes a segmentation foundation model, and we currently employ SAM (Segment Anything Model) [kirillov2023segment] for this role. SAM works in a class-agnostic manner and can generate high-quality segmentation results for the input image solely based on spatial prompt conditions. Specifically, for each class-specific prompt $\mathbf{p}_{k}$ (from above section), SAM outputs a corresponding mask $\mathcal{M}_{k}$ and a confidence score. These individual class-specific masks are aggregated into a final semantic segmentation map.
It is worth noting that the CAS module is frozen in the whole continual learning process. However, when the model learns task $t$, the outputs from CAS are used to generate the segmentation loss function. In our framework, the same segmentation loss as that for SAM training is adopted to train the learnable class-specific prompt generators, and the loss is also used together with the asymmetric loss [ridnik2021asymmetric] to train the task-specific LoRAs in the LTCD module.
During model inference, given a test image, the task-specific and class-specific components of each task are plugged into the segmentation system to segment image regions specifically for those classes estimated appearing in the input image. Such inference process is run over all tasks and the results are collected for the final segmentation. The final segmentation is created by aggregating individual class masks. While pixel ownership is unambiguous for the vast majority of pixels, any rare spatial overlaps are resolved using a confidence-based approach. The pixel is assigned to the class with the highest confidence score from the SAM decoder. Notably, this process requires no task ID, with time linear complexity with respect to the number of learned tasks $T$. Since only one adapter resides in memory at any time, maximum memory usage remains identical to that in single-task models. For time-sensitive applications, memory-time tradeoff techniques [s-lora] can be used to reduce inference latency.
@@
Data-free MiB* [CVPR2020ModelingTB] 39.02 16.73 31.29 36.68 9.81 27.77 34.22 5.26 24.29
PLOP* [douillardPLOPLearningForgetting2021] 40.38 13.41 31.52 39.46 12.58 30.10 38.12 7.32 27.39
@@ 4.1 Experimental Setup
Datasets Following previous studies [CVPR2020ModelingTB, douillardPLOPLearningForgetting2021], we evaluate our method on PASCAL VOC2012 [everinghamPascalVisualObject2010] and ADE20K [zhouSceneParsingADE20K2017]. PASCAL VOC2012 contains 20 object classes while ADE20K presents a more challenging scenario with 150 classes. Dataset statistics are in Appendix C.
CSS Settings. The training process is divided into $T$ tasks with certain protocol. The overlapped protocol, which is more challenging and realistic compared to the disjoint setting, allows images to contain old classes previously learned and future classes to be learned [SATS]. For these images, annotations are provided only for the current task’s classes and image regions corresponding to both old and future classes are annotated as background. Our experiments use this overlapped setting throughout. For instance, in the VOC2012 10-1 (11 tasks) setting, the model first learns to segment 10 classes, then incrementally learns one new class in each of 10 new tasks.
Implementation details. The model was trained for 5 epochs on PASCAL VOC2012 and 20 epochs on ADE20K using AdamW [loshchilov2017decoupled] optimizer (initial learning rate $1\times 10^{-4}$, weight decay 0.05) with a polynomial learning rate schedule. For the LTCD module, we set the similarity threshold $\tau$ to 0.3 for PASCAL VOC2012 and 0.2 for ADE20K, ensuring an appropriate quantity of selected visual tokens in $\mathbf{V}_{i}^{\text{sel}}$, with $M=30$ descriptive phrases per class and LoRA rank $L=32$. The SPG module uses $m=6$ prompt tokens with a single hidden layer, and pGen input lengths of 512 and 1024 for PASCAL VOC2012 and ADE20K respectively, with the longer length for ADE20K accommodating its complex scene composition and higher object density. All experiments were conducted with batch size 24 on 8 NVIDIA 3090 GPUs.
@@ 4.2 Main Results
MiB [CVPR2020ModelingTB] 3.27 6.35 5.91 7.08 6.72 6.80 39.50 20.07 24.70
PLOP [douillardPLOPLearningForgetting2021] 24.96 5.38 8.18 15.41 6.55 8.67 28.54 15.95 18.95
MiB+NeST [ECCV2024EarlyPP] 13.44 6.72 7.68 10.09 7.51 7.77 38.31 21.3 25.35
PLOP+NeST [ECCV2024EarlyPP] 27.98 6.53 8.26 17.57 7.84 10.15 33.89 17.84 21.66
@@ 4.5 Additional Studies
Finally, we analyse the computational cost of our approach. Similar to previous continual learning studies which employ task-specific modules [Li_2025_CVPR], our method introduces computational overhead during inference primarily from switching task-specific LoRA parameters. The inference time for a single task is 0.3 second per image, and the full process for six tasks in the (15-1) setting takes about 1.9 seconds per image. In comparison, most existing single encoder-decoder based CSS methods takes around 0.5 second per image. Despite leveraging foundation models, our parameter growth remains highly manageable. Each class requires only 8MB for the pGen module (0.225% of total model size), while all tasks’ LoRA adapters collectively need merely 22.18MB (0.62% of total model size). Each task requires relatively few trainable parameters per increment [ji2023continual], yet achieves state-of-the-art performance with remarkable gains in challenging settings, demonstrating practical applicability comparable to offline-trained models. Given our method’s substantial performance gains and affordable storage overhead, this trade-off is well-justified for non-realtime applications.